Yang Co. is expected to pay a $21.00 dividend next year. The dividend will decline by 10 percent
annually for the following three years. In year 5, Yang will sell off assets worth $100 per share. The
year 5 dividend, which includes a distribution of some of the proceeds of the asset sale, is expected to
be $60. In year 6, the dividend is expected to decrease to $40 and will be maintained at $40 for one
additional year. The dividend is then expected to grow by 5 percent annually thereafter. If the required
rate of return is 12 percent, what is the value of one share of Yang?
Answer #1

Annual dividends are

Year Dividend
1 21
2 21*90%=18.9
3 18.9*90%= 17.01
4 17.01*90%= 15.309
5 60
6 40
7 40

Horizon value= P7=D8/(k-g)

= 40*105%/(12%-5%) = 600

Current price = D1/(1+k)+ D2/(1+k)^2 +... (Dn+Pn) /(1+k)^n

= 21/1.12^1 +18.9/1.12^2+17.01/1.12^3+15.309/1.12^4+60/1.12^5+40/1.12^6+(40+600)/1.12^7

= $399.47